# Model Poisoning Attacks to Federated Learning via Multi-Round Consistency

### v2

@@ 5.2 Formulating an Optimization Problem
All-round optimization problem:  Suppose we have $n$ genuine clients and $m$ fake clients. The genuine clients are numbered as $1,2,\cdots,n$, while the fake clients are numbered as $n+1,\cdots,n+m$. We denote by $T$ the total number of training rounds and $\bm{w}^{T}$ the final global model. Recall that our attack goal is to poison the training process such that the final global model $\bm{w}^{T}$ has a large testing error rate. We observe that $\bm{w}^{T}=\bm{w}^{0}+\sum_{t=1}^{T}\bm{g}^{t}$, where $\bm{w}^{0}$ is the initial global model and $\bm{g}^{t}=AR(\{\bm{g}_{i}^{t}\}_{i\in[1,n+m]})$ is the aggregated model update in training round $t$. We denote by $\text{sign}(\sum_{t=1}^{T}\bm{g}^{t})$ the sign vector (i.e., update direction) of the total aggregated model update $\sum_{t=1}^{T}\bm{g}^{t}$ in the $T$ training rounds. Each dimension of a sign vector is either +1 or -1, which means increasing or decreasing the corresponding parameter of the global model.
If the total aggregated model update $\sum_{t=1}^{T}\bm{g}^{t}$ has a large magnitude and a random update direction, then the final global model $\bm{w}^{T}$ is substantially moved along the random update direction and would have a large testing error rate no matter what the initial global model $\bm{w}^{0}$ is. For instance, Fig. 10 in Appendix shows an example where a global model becomes nearly random guessing when moved along a random update direction substantially. Based on this observation, our goal is to craft the malicious model updates $\{\bm{g}_{i}^{t}\}_{i\in[n+1,n+m],\ t\in[1,T]}$ on the fake clients in the $T$ training rounds such that the total aggregated model update $\sum_{t=1}^{T}\bm{g}^{t}$ has a large magnitude along a random update direction. Formally, we formulate the following all-round optimization problem:
@@
[…]
@@ 5.3 Solving the Optimization Problem
It is still challenging to solve the per-round optimization problem in Equation 3. This is because $\bm{g}^{t}=AR(\{\bm{g}_{i}^{t}\}_{i\in[1,n+m]})$, but the attacker does not have access to the aggregation rule $AR$ nor the model updates of the genuine clients. We address the challenge via proposing a method to approximately solve the optimization problem. Our key idea is to enforce that the malicious model updates on the fake clients have the same sign vector $\bm{s}$, so the aggregated model updates are likely to have the sign vector $\bm{s}$ (see Table 4 in Appendix for an example in our experiments). Specifically, we assume that the malicious model update $\bm{g}_{i}^{t}$ of fake client $i$ in training round $t$ has the following form:
@@
Dynamically adjusting the unit magnitude vector $\bm{v}^{t}$:  A straightforward approach to pick the unit magnitude vector $\bm{v}^{t}$ is to set each dimension to the same value, which means that we treat all dimensions equally. However, as our experiments in Section 6 show, this strategy achieves suboptimal attack effectiveness. This is because the dimensions of the model updates on the genuine clients are unlikely to have the same magnitude. As a result, the model updates of the fake clients are dissimilar to those of the genuine clients, making them filtered out when a defense is deployed on the server. It is challenging to construct a unit magnitude vector similar to those of the genuine model updates, since the attacker does not have access to the model updates on the genuine clients in our threat model.
To address this challenge, we observe that the global model difference $\bm{w}^{t-1}-\bm{w}^{t-2}$ is the aggregated model update $\bm{g}^{t-1}$ in the previous training round $t-1$. Moreover, $\bm{g}^{t-1}$ is an aggregation of the model updates on the genuine clients and those on the fake clients. The attacker knows $\bm{w}^{t-1}$, $\bm{w}^{t-2}$, and the model updates $\bm{g}_{i}^{t-1}=\bm{k}^{t-1}\odot\bm{s}$ on the fake clients in the previous training round. Therefore, the attacker can estimate the unit magnitude vector of the genuine clients in the previous training round and use it as $\bm{v}^{t}$ in the current training round. In particular, we can normalize the model update $\bm{k}^{t-1}\odot\bm{s}$ on the fake clients in the previous training round to have the same magnitude as $\bm{g}^{t-1}$ and then subtract it from $\bm{g}^{t-1}$ to estimate the model updates on the genuine clients. Formally, we have the following:
@@
where $\bm{g}^{t-1}=\bm{w}^{t-1}-\bm{w}^{t-2}$ is the aggregated model update in the previous training round, $\frac{\left\|\bm{w}^{t-1}-\bm{w}^{t-2}\right\|}{\left\|\bm{k}^{t-1}\odot\bm{s}\right\|}$ is used to normalize $\bm{k}^{t-1}\odot\bm{s}$ to have the same magnitude as $\bm{g}^{t-1}$ so they are on the same scale for subtraction, the absolute value ||$\cdot$|| ⋅ | is used in the numerator because $\bm{v}^{t}$ is a magnitude vector, and the denominator is used to normalize $\bm{v}^{t}$ to have $\ell_{2}$-norm of 1.
Dynamically adjusting the scaling factor $\lambda^{t}$:  A naive approach to pick $\lambda^{t}$ is to set it as a large number, which means using model updates with large magnitudes on the fake clients. However, as our experiments in Section 6 show, this approach achieves suboptimal attack effectiveness because the malicious model updates may be filtered out by a defense. To address the challenge, we dynamically set $\lambda^{t}$ based on the global model difference (i.e., aggregated model update) in the previous training round. Our intuition is that, if the aggregated model update in the previous training round has a large (or small) magnitude, the server may also expect an aggregated model update with a large (or small) magnitude in the current training round. Based on this intuition, we set $\lambda^{t}$ to be proportional to the magnitude of the aggregated model update (i.e., $\left\|\bm{w}^{t-1}-\bm{w}^{t-2}\right\|$) in the previous training round. Formally, we have the following:
@@
Dynamically adjusting $c^{t}$ based on hypothesis testing:  We adjust $c^{t}$ in training round $t$ based on whether the attack is successful in the past $e$ training rounds. Specifically, we formulate the following two hypothesis:
@@
If the hypothesis testing favors $H_{1}$ with a p-value $p$, $c^{t}$ does not change, i.e., we set $c^{t}=c^{t-1}$; otherwise the malicious model updates may have been filtered out by the defense in the past training rounds and thus we decrease $c^{t}$ by a factor, i.e., $c^{t}=\beta c^{t-1}$, where $\beta<1$. Specifically, the attacker can calculate the total aggregated model update in the past $e$ training rounds, i.e., $\bm{w}^{t-1}-\bm{w}^{t-e}$. Moreover, the attacker can calculate the number (denoted as $X$) of dimensions of $\bm{w}^{t-1}-\bm{w}^{t-e}$ whose signs match with those in $\bm{s}$. Under the null hypothesis $H_{0}$, the malicious model updates were filtered out by the defense, and thus the signs of the total aggregated model update $\bm{w}^{t-1}-\bm{w}^{t-e}$ are not correlated with $\bm{s}$. Since $\bm{s}$ is picked uniformly at random, $X$ is a random variable and follows a binomial distribution $Bin(d,0.5)$, where $d$ is the number of dimensions of the global model. Therefore, if we have $\text{Pr}(x\geq X)\leq p$, then we reject $H_{0}$ and favor $H_{1}$, where $x\sim Bin(d,0.5)$ and $p$ is the p-value. We use $p=0.01$ in experiments.
6 Evaluation
6.1 Experimental Setup
6.1.1 Datasets
We use the following five datasets from different domains.
6.1.2 Defenses and Compared Attacks
We evaluate eight state-of-the-art defenses, including six Byzantine-robust aggregation rules, i.e., Multi-Krum (M-Krum) [12], Median [17], Trimmed Mean (TrMean) [17], Norm Bound (N-Bound) [16], FLTrust [13], and FLAME [15], one provably robust defense, i.e., FLCert [14], and one malicious clients detection method, i.e., FLDetector (FLDet) [18]. We also consider the non-robust FedAvg [2] as a baseline. We compare our PoisonedFL with seven attacks, including five attacks (Fang [10], Opt. Fang [11], LIE [8], Min-Max [11], and Min-Sum [11]) that require genuine clients’ information and two attacks (Random [9] and MPAF [9]) that do not require such information. Details are deferred to Appendix B.2 and B.3.
6.1.3 Evaluation Metric
6.1.4 FL Settings
By default, we assume $1,200$ genuine clients and 20% of fake clients, i.e., the number of fake clients is 20% of the number of genuine clients. In each training round, the server selects a random subset of 10% of clients to participate in training. A genuine client trains its local model for 1 local epoch in a training round using SGD. Due to limited space, more FL settings for the five datasets are detailed in Appendix B.1. We also study the impact of different FL settings (fraction of fake clients, non-IID degree, local training epochs, and participation rate) in Section 6.3. Our PoisonedFL has parameters $c^{t}$, $e$, and $\beta$. We set $c^{0}=8$, $e=50$, and $\beta=0.7$, and we do not decrease $c^{t}$ to be smaller than $0.5$ since a too small $c^{t}$ is ineffective.
@@ 6.1.4 FL Settings
6.2 Main Results
Table 1 shows the testing error rates of the learnt global models under different attacks and defenses for the five datasets. “No Attack” means no fake clients are injected. We have the following observations.
PoisonedFL breaks state-of-the-art FL defenses:  PoisonedFL substantially increases the testing error rates of the global models learnt by FedAvg and state-of-the-art defenses, compared to No Attack. Specifically, PoisonedFL increases the testing error rate by 15.84 to 90.67 across the five datasets and nine FL methods. Moreover, in 29 out of the 45 cases, PoisonedFL makes the global model nearly random guessing or worse. Our results show that existing FL defenses are not robust even if an attacker has no knowledge about the defense nor the genuine clients’ local data or models.
Why PoisonedFL breaks FL defenses:  State-of-the-art defenses leverage filtering and clipping to reduce the impact of malicious model updates. They can weaken the attack effect in individual training rounds. However, since PoisonedFL crafts malicious model updates that are consistent across training rounds and avoid being entirely filtered out by dynamically adjusting attack magnitudes, its cumulative attack effect still substantially moves the global model towards the designated random update direction $\bm{s}$, leading to a large testing error rate. For instance, Table 4 in Appendix shows that a large fraction (over 90% percent) of the dimensions of the total aggregated model updates have signs matching with $\bm{s}$ and Table 5 in Appendix shows that the magnitude of the total aggregated model updates becomes very large, after our attack effect accumulates over multiple training rounds, for all defenses. The detailed analysis for each defense is defered to Appendix A.1.
PoisonedFL outperforms existing attacks:  For the non-robust FedAvg, both PoisonedFL and some existing attacks can reduce the global models to random guessing or worse. However, for the five datasets and eight FL defenses, our PoisonedFL achieves a much larger testing error rate than existing attacks in 32 out of the 40 cases with the gap ranging from 2.11 to 84.33. In six of the remaining cases, both our PoisonedFL and some existing attacks make the learnt global models nearly random guessing or worse. Our PoisonedFL achieves a smaller testing error rate than some existing attacks only in two cases. In particular, Fang achieves a testing error rate of 70.02 for the case (Purchase, TrMean), while PoisonedFL achieves a testing error rate of 53.86 in the case; and Min-Max (or Min-Sum) achieves a testing error rate of 94.88 (or 94.97) for (FEMNIST, FLDetector), while PoisonedFL achieves a testing error rate of 91.25 in the case. However, we note that PoisonedFL still achieves large testing error rates even in these rare cases.
Among the three attacks Random, MPAF, and PoisonedFL, which do not require knowledge about the local training data/models on the genuine clients nor the defense, Random is ineffective in most cases when a defense is deployed; while MPAF is effective to some extent at attacking some defenses, e.g., TrMean and Norm Bound, but ineffective at attacking other defenses, e.g., Multi-Krum, FLTrust, and FLDetector. PoisonedFL substantially outperforms Random and MPAF when a defense is deployed in all cases except (CIFAR10, Norm Bound), for which both MPAF and PoisonedFL make the global models random guessing. For instance, on MNIST dataset, PoisonedFL achieves 59.24 to 85.87 higher testing error rates than MPAF at attacking a defense.
Why PoisonedFL outperforms existing attacks:  PoisonedFL outperforms existing attacks because their crafted malicious model updates are inconsistent across training rounds, leading to self-cancellation of attack effect. In contrast, PoisonedFL enforces the malicious model updates to be consistent (i.e., have the same sign vector) across training rounds, making the aggregated model updates add up in the global model across training rounds. As a result, the final learnt global model has a very large magnitude, leading to a large testing error rate. For instance, Table 5 in Appendix shows the magnitude (measured by $\ell_{2}$ norm) of the total aggregated model updates for different attacks on the Purchase dataset. Our results show that the total aggregated model updates under existing attacks have small magnitudes while those under our PoisonedFL attack have much larger magnitudes, which is caused by the inconsistent malicious model updates in existing attacks vs. consistent malicious model updates in our attack.
6.3 Impact of FL Settings
We study the impact of different FL settings, including the fraction of fake clients, degree of non-IID, the number of local training epochs, and fraction of clients that are selected to participate in each training round, on the effectiveness of existing attacks and our attack. Unless otherwise mentioned, we show experimental results on CIFAR-10 dataset for simplicity, but we have similar observations on other datasets. Moreover, when studying the impact of one parameter, we use the default settings for other parameters discussed in Section 6.1.4.
(a) M-Krum
@@ 6.3 Impact of FL Settings
(h) FLDet
Figure 3: Testing error rate of the global model as a function of the fraction of fake clients (%) under different defenses and attacks.
(a) M-Krum

### v3

@@ 4.2 Formulating an Optimization Problem
All-round optimization problem:  Suppose we have $n$ genuine clients and $m$ fake clients. The genuine clients are numbered as $1,2,\cdots,n$, while the fake clients are numbered as $n+1,\cdots,n+m$. We denote by $T$ the total number of training rounds and $\bm{w}^{T}$ the final global model. Recall that our attack goal is to poison the training process such that the final global model $\bm{w}^{T}$ has a large testing error rate. We observe that $\bm{w}^{T}=\bm{w}^{0}+\sum_{t=1}^{T}\bm{g}^{t}$, where $\bm{w}^{0}$ is the initial global model and $\bm{g}^{t}=AR(\{\bm{g}_{i}^{t}\}_{i\in[1,n+m]})$ is the aggregated model update in training round $t$. We denote by $\text{sign}(\sum_{t=1}^{T}\bm{g}^{t})$ the sign vector (i.e., update direction) of the total aggregated model update $\sum_{t=1}^{T}\bm{g}^{t}$ in the $T$ rounds.
If the total aggregated model update $\sum_{t=1}^{T}\bm{g}^{t}$ has a large magnitude and a random update direction, then the final global model $\bm{w}^{T}$ is substantially moved along the random update direction and would have a large testing error rate no matter what the initial global model $\bm{w}^{0}$ is. For instance, Fig. 6 in Appendix shows an example where a global model becomes nearly random guessing when moved along a random update direction substantially. Based on this observation, our goal is to craft the malicious model updates $\{\bm{g}_{i}^{t}\}_{i\in[n+1,n+m],\ t\in[1,T]}$ on the fake clients in the $T$ rounds such that the total aggregated model update $\sum_{t=1}^{T}\bm{g}^{t}$ has a large magnitude along a random update direction. Formally, we formulate the following all-round optimization problem:
@@
where $\left\|\cdot\right\|$ is $\ell_{2}$-norm and $\bm{s}$ is a random sign vector that characterizes a random update direction. Each dimension of $\bm{s}$ is picked as either +1 or -1 uniformly at random. However, the all-round optimization problem is hard to solve. In particular, when solving the malicious model updates in training round $t$, the model updates in the future training rounds are not available yet. Therefore, we transform our all-round optimization problem into a per-round one.
@@ 4.3 Solving the Optimization Problem
It is still challenging to solve the per-round optimization problem in Equation 3. This is because $\bm{g}^{t}=AR(\{\bm{g}_{i}^{t}\}_{i\in[1,n+m]})$, but the attacker does not have access to the aggregation rule $AR$ nor the model updates of the genuine clients. We address the challenge via proposing a method to approximately solve the optimization problem. Our key idea is to enforce that the malicious model updates on the fake clients have the same sign vector $\bm{s}$, so the aggregated model updates are likely to have the sign vector $\bm{s}$ (see Fig. 7 in Appendix for an example in our experiments). Specifically, we assume that the malicious model update $\bm{g}_{i}^{t}$ of fake client $i$ in training round $t$ has the following form:
@@
Dynamically adjusting the unit magnitude vector $\bm{v}^{t}$:  A straightforward approach to pick the unit magnitude vector $\bm{v}^{t}$ is to set each dimension to the same value, which means that we treat all dimensions equally. However, as our experiments in Section 5 show, this strategy achieves suboptimal attack effectiveness. This is because the dimensions of the model updates on the genuine clients are unlikely to have the same magnitude. As a result, the model updates of the fake clients are dissimilar to those of the genuine clients, making them be filtered out when a defense is deployed on the server. It is challenging to construct an unit magnitude vector similar to those of the model updates on the genuine clients, since the attacker does not have access to the model updates on the genuine clients in our threat model.
To address this challenge, we observe that the global model difference $\bm{w}^{t-1}-\bm{w}^{t-2}$ is the aggregated model update $\bm{g}^{t-1}$ in the previous training round t1t-1. Moreover, $\bm{g}^{t-1}$ is an aggregation of the model updates on the genuine clients and those on the fake clients. The attacker knows $\bm{w}^{t-1}$, $\bm{w}^{t-2}$, and the model updates $\bm{g}_{i}^{t-1}=\bm{k}^{t-1}\odot\bm{s}$ on the fake clients in the previous training round. Therefore, the attacker can estimate the unit magnitude vector of the genuine clients in the previous training round and use it as $\bm{v}^{t}$ in the current training round. In particular, we can normalize the model update $\bm{k}^{t-1}\odot\bm{s}$ on the fake clients in the previous training round to have the same magnitude as $\bm{g}^{t-1}$ and then subtract it from $\bm{g}^{t-1}$ to estimate the model updates on the genuine clients. Formally, we have the following:
@@
where $\bm{g}^{t-1}=\bm{w}^{t-1}-\bm{w}^{t-2}$ is the aggregated model update in the previous training round, $\frac{\left\|\bm{w}^{t-1}-\bm{w}^{t-2}\right\|}{\left\|\bm{k}^{t-1}\odot\bm{s}\right\|}$ is used to normalize $\bm{k}^{t-1}\odot\bm{s}$ to have the same magnitude as $\bm{g}^{t-1}$ so they are on the same scale for subtraction, the absolute value $|\cdot|$ is used in the numerator because $\bm{v}^{t}$ is a magnitude vector, and the denominator is used to normalize $\bm{v}^{t}$ to have $\ell_{2}$-norm of 1.
Dynamically adjusting the scaling factor $\lambda^{t}$:  A naive approach to pick $\lambda^{t}$ is to set it as a large number. However, as our experiments in Appendix C.3 show, this approach may be filtered out by a defense. To address the challenge, we dynamically set $\lambda^{t}$ based on the global model difference (i.e., aggregated model update) in the previous training round. Our intuition is that, if the aggregated model update in the previous training round has a large (or small) magnitude, the server may also expect an aggregated model update with a large (or small) magnitude in the current training round. Therefore, we set $\lambda^{t}$ to be proportional to the magnitude of the aggregated model update (i.e., $\left\|\bm{w}^{t-1}-\bm{w}^{t-2}\right\|$) in the previous training round. Formally, we have the following:
@@
Dynamically adjusting $c^{t}$ based on hypothesis testing:  We adjust $c^{t}$ in training round $t$ based on whether the attack is successful in the past $e$ training rounds. Specifically, we formulate the following two hypotheses:
@@
If the hypothesis testing favors $H_{1}$ with a p-value $p$, $c^{t}$ does not change, i.e., we set $c^{t}=c^{t-1}$; otherwise the malicious model updates may have been filtered out by the defense in the past training rounds and thus we decrease $c^{t}$ by a factor, i.e., $c^{t}=\beta c^{t-1}$, where $\beta<1$. Specifically, the attacker can calculate the total aggregated model update in the past $e$ training rounds, i.e., $\bm{w}^{t-1}-\bm{w}^{t-e}$. Moreover, the attacker can calculate the number (denoted as $X$) of dimensions of $\bm{w}^{t-1}-\bm{w}^{t-e}$ whose signs match with those in $\bm{s}$. Under the null hypothesis $H_{0}$, the malicious model updates were filtered out by the defense and thus the signs of the total aggregated model update $\bm{w}^{t-1}-\bm{w}^{t-e}$ are not correlated with $\bm{s}$. Since $\bm{s}$ is picked uniformly at random, $X$ is a random variable and follows a binomial distribution $Bin(d,0.5)$, where $d$ is the number of dimensions of the global model. Therefore, if we have $\text{Pr}(x\geq X)\leq p$, then we reject $H_{0}$ and favor $H_{1}$, where $x\sim Bin(d,0.5)$ and $p$ is the p-value. We use $p=0.01$ in experiments.
@@
5 Evaluation
5.1 Experimental Setup
5.1.1 Datasets
We use five datasets from different domains, including MNIST [24], FashionMNIST [35], Purchase [7], CIFAR-10 [23], and FEMNIST [12]. Detailed introduction and default non-IID degrees are deferred to Appendix B.1.
5.1.2 Evaluation Metric
5.1.3 Defenses and Compared Attacks
We evaluate eight state-of-the-art defenses, including six Byzantine-robust aggregation rules (Multi-Krum [11], Median [39], Trimmed Mean [39], Norm Bound [34], FLTrust [14], and FLAME [28]), one provably robust defense (FLCert [16]) and one malicious clients detection method (FLDetector [40]). We also consider the non-robust FedAvg [26] as a baseline. We compare our PoisonedFL with seven attacks, including five attacks (Fang [17], Opt. Fang [32], LIE [10], Min-Max [32], and Min-Sum [32]) that require genuine clients’ information and two attacks (Random [13] and MPAF [13]) that do not require such information. Detailed introduction is deferred to Appendix B.2 and B.3.
(a) Multi-Krum
@@ 5.1.3 Defenses and Compared Attacks
(h) FLDetector
Figure 3: Testing error rate of the global model as a function of the fraction of fake clients (%) under different scenarios.
(a) Multi-Krum